OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? show your work plz

OpenStudy (anonymous):

x + y = 150 25x + 28.5y = 4770 Can you solve this system of equations?

OpenStudy (anonymous):

A=Adults S=Students A45+S24.5=4770 A+S=150

OpenStudy (anonymous):

Hint: Use substitution